OpenStudy (anonymous):

Express the equation in Slope-intercept form: 10x+3y=30

OpenStudy (anonymous):

@kyria the slope intercept form is y =mx+c soo we need to divide the eqn 10 x+3y=30 by 3 oo we get 10/3x+y=10 arrenging it we get y =-10/3x +10 .